An electric flied is given by x^3kN/C.find the potential difference between the point on the x axis at x=1m and x=2m

1
Expert's answer
2022-06-06T08:54:43-0400

"\varphi=\int Edx=\int_1^2 x^3dx=\frac{x^4}4|_1^2=3.75~kV."